OpenStudy (maheshmeghwal9):

How can u prove that sqrrt{2} is Irrational?

OpenStudy (henryblah):

I can't explain it well. Perhaps this page will help. http://www.homeschoolmath.net/teaching/proof_square_root_2_irrational.php

Parth (parthkohli):

You can't express it in the form of \(\Large \color{purple}{p \over q}\) Can you? No, you can't. p over q actually means that you express it in a form of a fraction. You can't in this case. You can actually, \(\Large {{sqrt{2} \over 1}}\), but this isn't really a fraction.

OpenStudy (lgbasallote):

the better term is "express as quotient of two integers"

OpenStudy (maheshmeghwal9):

But I don't get it really what you are saying.Please tell me in more easy way.

OpenStudy (lgbasallote):

@ParthKohli

Parth (parthkohli):

I was about to type that.

OpenStudy (maheshmeghwal9):

Ya ok!

OpenStudy (mertsj):

Assume that it is rational and then show that it cannot be expressed as the quotient of two integers.

OpenStudy (anonymous):

If you are geometrically minded you might like this one http://jeremykun.wordpress.com/2011/08/14/the-square-root-of-2-is-irrational-geometric-proof/

OpenStudy (anonymous):

right ... Let's suppose √2 were a rational number. Then we can write it √2 = a/b where a, b are whole numbers, b not zero. We additionally make it so that this a/b is simplified to the lowest terms, since that can obviously be done with any fraction. It follows that 2 = a2/b2, or a2 = 2 * b2. So the square of a is an even number since it is two times something. From this we can know that a itself is also an even number. Why? Because it can't be odd; if a itself was odd, then a * a would be odd too. Odd number times odd number is always odd. Check if you don't believe that! Okay, if a itself is an even number, then a is 2 times some other whole number, or a = 2k where k is this other number. We don't need to know exactly what k is; it won't matter. Soon is coming the contradiction: If we substitute a = 2k into the original equation 2 = a2/b2, this is what we get:2 = (2k)2/b2 2 = 4k2/b2 2*b2 = 4k2 b2 = 2k2. This means b2 is even, from which follows again that b itself is an even number!!! WHY is that a contradiction? Because we started the whole process saying that a/b is simplified to the lowest terms, and now it turns out that a and b would both be even. So √2 cannot be rational.

OpenStudy (maheshmeghwal9):

But how mertsj! please tell me.

OpenStudy (lgbasallote):

lol it's always fun when people suddenly post more complicated ways of solving the problem >:))) hahaha

Parth (parthkohli):

@badi Lol, integers. You can put negative values too in the \(\Large {p \over q}\) thingy.

Parth (parthkohli):

@lgbasallote hahahah

OpenStudy (maheshmeghwal9):

lol<:((((((((((((((((((((((((((((((((((((((((

OpenStudy (maheshmeghwal9):

Thanx! Badi

Parth (parthkohli):

@maheshmeghwal9 don't complicate things in your mind.

OpenStudy (anonymous):

this sol btw is copy-pasted from some website right :)

Parth (parthkohli):

http://www.homeschoolmath.net/teaching/proof_square_root_2_irrational.php COPIED FROM HERE

OpenStudy (mertsj):

Can the asker do the problem now? I know that too many cooks spoil the broth.

OpenStudy (anonymous):

i think i did this last year though i dont remember it now u had to assume rt2 to be a rational no and prove that p and q were not co-prime or something

OpenStudy (maheshmeghwal9):

Ya! dude

OpenStudy (anonymous):

many hands make light work....

OpenStudy (maheshmeghwal9):

I mean Mertsj!

OpenStudy (mertsj):

@estudier That may be true for jobs that actually involve hands. With minds...not so much.

OpenStudy (maheshmeghwal9):

oh

Parth (parthkohli):

Opposing proverbs @estudier @Mertsj

OpenStudy (anonymous):

I don't know, community math (and other) solving seems to be growing in popularity. Think it depends more on the nature of the problem and how good the group facilities are

OpenStudy (lgbasallote):

let me share one too "if you want to live a happy life, tie it to a goal, not a person or objects ~Einstein"

OpenStudy (vincent-lyon.fr):

Here is badi's solution sung by a German pupil in secondary school: http://www.youtube.com/watch?v=vnMrUAvKtAg Hope you enjoy it, even if you do not understand the language!